OpenStudy (anonymous):

a bullet leaves a rifle with a muzzle velocity of 521 m/s. while accelerating through the barrel of the rifle, the bullet moves a distance of 0.840 m. determine the acceleration of the bullet. ( assume uniform acceleration)

OpenStudy (anonymous):

\[v ^{2}=u ^{2}+2as\] \[521^{2}=0+ 2\times a \times 0.840\] \[a= 161572.0238m/s ^{2}\]